find the value of sin4A and cos4A
Answer by fractalier(6550)

Just posted sin 4a...but
cos 4a = cos 2(2a) =
cos^2 2a - sin^2 2a =
(cos^2 a - sin^2 a)^2 - (2 sin a cos a)^2 =
cos^4 a - 2 sin^2 a cos^2 a + sin^4 a - 4 sin^2 a cos^2 a =
cos^4 a + sin^4 a - 6 sin^2 a cos^2 a
Further manipulation is possible.